\begin{thebibliography}{}

\bibitem[\protect\citename{Bojar \bgroup et al.\egroup
  }2014]{bojar:2014:monohindi}
Ond{\v r}ej Bojar, Vojt{\v e}ch Diatka, Pavel Rychl{\'y}, Pavel Stra{\v
  n}{\'a}k, V{\'i}t Suchomel, Ale{\v s} Tamchyna, and Daniel Zeman.
\newblock 2014.
\newblock {HindMonoCorp} 0.5.
\newblock \url{http://hdl.handle.net/11858/00-097C-0000-0023-6260-A}.
\newblock {LINDAT}/{CLARIN} digital library at the Institute of Formal and
  Applied Linguistics, Charles University in Prague.

\bibitem[\protect\citename{Brown \bgroup et al.\egroup }1990]{brown:1990}
Peter~F Brown, John Cocke, Stephen A~Della Pietra, Vincent J~Della Pietra,
  Fredrick Jelinek, John~D Lafferty, Robert~L Mercer, and Paul~S Roossin.
\newblock 1990.
\newblock {A Statistical Approach to Machine Translation}.
\newblock {\em Computational linguistics}, 16(2):79--85.

\bibitem[\protect\citename{Chen and Goodman}1996]{chen:1996}
Stanley~F Chen and Joshua Goodman.
\newblock 1996.
\newblock {An Empirical Study of Smoothing Techniques for Language Modeling}.
\newblock In {\em Proceedings of the 34th annual meeting on Association for
  Computational Linguistics}, pages 310--318. Association for Computational
  Linguistics.

\bibitem[\protect\citename{Doddington}2002]{doddington:2002}
George Doddington.
\newblock 2002.
\newblock {Automatic Evaluation of Machine Translation Quality using N-gram
  Co-occurrence Statistics}.
\newblock In {\em Proceedings of the second international conference on Human
  Language Technology Research}, pages 138--145. Morgan Kaufmann Publishers
  Inc.

\bibitem[\protect\citename{Durrani \bgroup et al.\egroup }2014]{durrani:2014}
Nadir Durrani, Hassan Sajjad, Hieu Hoang, and Philipp Koehn.
\newblock 2014.
\newblock Integrating an unsupervised transliteration model into statistical
  machine translation.
\newblock In {\em Proceedings of the 15th Conference of the European Chapter of
  the ACL (EACL 2014)}, volume~14, pages 148--153. Association for
  Computational Linguistics.

\bibitem[\protect\citename{Heafield}2011]{heafield:2011}
Kenneth Heafield.
\newblock 2011.
\newblock {KenLM: Faster and Smaller Language Model Queries}.
\newblock In {\em Proceedings of the Sixth Workshop on Statistical Machine
  Translation}, pages 187--197. Association for Computational Linguistics.

\bibitem[\protect\citename{Khapra \bgroup et al.\egroup }2010]{khapra:2010}
Mitesh~M Khapra, Anup Kulkarni, Saurabh Sohoney, and Pushpak Bhattacharyya.
\newblock 2010.
\newblock {All Words Domain Adapted WSD: Finding a Middle Ground between
  Supervision and Unsupervision}.
\newblock In {\em Proceedings of the 48th Annual Meeting of the Association for
  Computational Linguistics}, pages 1532--1541. Association for Computational
  Linguistics.

\bibitem[\protect\citename{Koehn and Hoang}2007]{koehn:2007:factored}
Philipp Koehn and Hieu Hoang.
\newblock 2007.
\newblock Factored translation models.
\newblock In {\em EMNLP-CoNLL}, pages 868--876.

\bibitem[\protect\citename{Koehn and Knight}2003]{koehn:2003:CS}
Philipp Koehn and Kevin Knight.
\newblock 2003.
\newblock {Empirical Methods for Compound Splitting}.
\newblock In {\em Proceedings of the tenth conference on European chapter of
  the Association for Computational Linguistics-Volume 1}, pages 187--193.
  Association for Computational Linguistics.

\bibitem[\protect\citename{Koehn \bgroup et al.\egroup }2003]{koehn:2003}
Philipp Koehn, Franz~Josef Och, and Daniel Marcu.
\newblock 2003.
\newblock {Statistical Phrase-Based Translation}.
\newblock In {\em Proceedings of the 2003 Conference of the North American
  Chapter of the Association for Computational Linguistics on Human Language
  Technology-Volume 1}, pages 48--54. Association for Computational
  Linguistics.

\bibitem[\protect\citename{Koehn \bgroup et al.\egroup }2007]{koehn:2007}
Philipp Koehn, Hieu Hoang, Alexandra Birch, Chris Callison-Burch, Marcello
  Federico, Nicola Bertoldi, Brooke Cowan, Wade Shen, Christine Moran, Richard
  Zens, et~al.
\newblock 2007.
\newblock {Moses: Open Source Toolkit for Statistical Machine Translation}.
\newblock In {\em Proceedings of the 45th annual meeting of the ACL on
  interactive poster and demonstration sessions}, pages 177--180. Association
  for Computational Linguistics.

\bibitem[\protect\citename{Marcu and Wong}2002]{marcu:2002}
Daniel Marcu and William Wong.
\newblock 2002.
\newblock {A Phrase-Based, Joint Probability Model for Statistical Machine
  Translation}.
\newblock In {\em Proceedings of the ACL-02 conference on Empirical methods in
  natural language processing-Volume 10}, pages 133--139. Association for
  Computational Linguistics.

\bibitem[\protect\citename{Minnen \bgroup et al.\egroup }2001]{minnen:2001}
Guido Minnen, John Carroll, and Darren Pearce.
\newblock 2001.
\newblock Applied morphological processing of english.
\newblock {\em Natural Language Engineering}, 7(03):207--223.

\bibitem[\protect\citename{Och and Ney}2003]{och:2003}
Franz~Josef Och and Hermann Ney.
\newblock 2003.
\newblock {A Systematic Comparison of Various Statistical Alignment Models}.
\newblock {\em Computational linguistics}, 29(1):19--51.

\bibitem[\protect\citename{Papineni \bgroup et al.\egroup }2002]{papineni:2002}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu.
\newblock 2002.
\newblock {BLEU: A Method for Automatic Evaluation of Machine Translation}.
\newblock In {\em Proceedings of the 40th annual meeting on association for
  computational linguistics}, pages 311--318. Association for Computational
  Linguistics.

\bibitem[\protect\citename{Patel \bgroup et al.\egroup }2013]{patel:2013}
Raj~Nath Patel, Rohit Gupta, Prakash~B Pimpale, and Sasikumar M.
\newblock 2013.
\newblock Reordering rules for english-hindi smt.
\newblock In {\em Proceedings of the Second Workshop on Hybrid Approaches to
  Translation}, pages 34--41. Association for Computational Linguistics.

\bibitem[\protect\citename{Patel \bgroup et al.\egroup }2014]{patel:2014}
Raj~Nath Patel, Prakash~B Pimpale, and Sasikumar M.
\newblock 2014.
\newblock {Statistical Machine Translation for Indian Languages: Mission
  Hindi}.
\newblock In {\em Proceedings of NLP Tools contest}. ICON-2014: 11th
  International Conference on Natural Language Processing.

\bibitem[\protect\citename{Pimpale \bgroup et al.\egroup }2014]{pimpale:2014}
Prakash~B Pimpale, Raj~Nath Patel, and Sasikumar M.
\newblock 2014.
\newblock {SMT from Agglutinative Languages: Use of Suffix Separation and Word
  Splitting}.
\newblock In {\em Proceedings of ICON-2014: 11th International Conference on
  Natural Language Processing}.

\bibitem[\protect\citename{Popovic and Ney}2004]{popovic:2004:word-stem}
Maja Popovic and Hermann Ney.
\newblock 2004.
\newblock {Towards the Use of Word Stems and Suffixes for Statistical Machine
  Translation.}
\newblock In {\em LREC}.

\bibitem[\protect\citename{Popovi{\'c} \bgroup et al.\egroup
  }2006]{popovic:2006:CW}
Maja Popovi{\'c}, Daniel Stein, and Hermann Ney.
\newblock 2006.
\newblock {Statistical Machine Translation of German Compound Words}.
\newblock In {\em Advances in Natural Language Processing}, pages 616--624.
  Springer.

\bibitem[\protect\citename{Ramanathan and Rao}2003]{ramanathan:2003:stem}
Ananthakrishnan Ramanathan and Durgesh~D Rao.
\newblock 2003.
\newblock {A Lightweight Stemmer for Hindi}.
\newblock In {\em The Proceedings of EACL}.

\bibitem[\protect\citename{Snover \bgroup et al.\egroup }2006]{snover:2006:ter}
Matthew Snover, Bonnie Dorr, Richard Schwartz, Linnea Micciulla, and John
  Makhoul.
\newblock 2006.
\newblock {A Study of Translation Edit Rate with Targeted Human Annotation}.
\newblock In {\em Proceedings of association for machine translation in the
  Americas}, volume 200.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{acl2016}
\usepackage{times}
\usepackage[hyphens]{url}
\usepackage{latexsym}
\usepackage{xspace}
\aclfinalcopy 
\pdfoutput=1

\newcommand{\bnhi}{\textit{bn-hi}\xspace}
\newcommand{\enhi}{\textit{en-hi}\xspace}
\newcommand{\tehi}{\textit{te-hi}\xspace}
\newcommand{\tahi}{\textit{ta-hi}\xspace}
\newcommand{\mrhi}{\textit{mr-hi}\xspace}

\title{Statistical Machine Translation for Indian Languages: Mission Hindi 2}

\author{Raj Nath Patel \\
	KBCS, CDAC Mumbai \\
	{\tt rajnathp@cdac.in} \\\And
	Prakash B. Pimpale \\
	KBCS, CDAC Mumbai \\
	{\tt prakash@cdac.in} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
This paper presents Centre for Development of Advanced Computing Mumbai's (CDACM) submission to NLP Tools Contest on Statistical Machine Translation in Indian Languages (ILSMT) 2015 (collocated with ICON 2015). The aim of the contest was to collectively explore the effectiveness of Statistical Machine Translation (SMT) while translating within Indian languages and between English and Indian languages. In this paper, we report our work on all five language pairs, namely Bengali-Hindi (\bnhi), Marathi-Hindi (\mrhi), Tamil-Hindi (\tahi), Telugu-Hindi (\tehi), and English-Hindi (\enhi) for Health, Tourism and General domains. We have used suffix separation, compound splitting and preordering prior to SMT training and testing.
\end{abstract}


\section{Introduction}
\label{sec:intro}
In this paper, we present our SMT experiments from Bengali, Marathi, Tamil, Telugu and English to Hindi. From the set of languages involved, Bengali, Hindi and Marathi belong to the Indo-Aryan family and Tamil and Telugu are from Dravidian language family. All languages except English, have the same flexibility towards word order, canonically following
the Subject-Object-Verb (SOV) structure.

With reference to morphology Bengali, Marathi, Tamil, and Telugu are more agglutinative compared to Hindi. It is known that SMT produces unknown words resulting in bad translation quality if morphological divergence between source and target languages is high. ~\newcite{koehn:2003:CS}, ~\newcite{popovic:2004:word-stem}, nd~\newcite{popovic:2006:CW} have demonstrated ways to handle this issue with morphological segmentation of words in the source sentences before training the SMT system. To handle this morphological difference we have used suffix separation and compound word splitting developed by~\newcite{pimpale:2014} and~\newcite{patel:2014}.

For English-Hindi SMT, we achieve better alignment using preordering~\cite{patel:2013} and stem as an alignment factor~\cite{koehn:2007:factored}.

All machine translation (MT) systems suffer from Out of Vocabulary (OOV) words. These OOV words are mostly named entities, technical terms and foreign words which can be translated using transliteration system. We used ~\newcite{durrani:2014}, which is a fully unsupervised approach for developing a transliteration system using parallel corpus meant for the SMT training. 

The rest of the paper is organized as follows. In section~\ref{sec:settings}, we discuss dataset and experimental setup. Section~\ref{sec:results} discusses experiments and results. Submitted systems to the shared task are described in section~\ref{sec:submit} followed by the conclusion and future work in section~\ref{sec:conclution}.

\section{Data-set and Experimental Setup}
\label{sec:settings}
In the following subsections, we describe Training and Testing corpus followed by pre-processing and SMT system setup for experiments.

\subsection{Corpus for SMT Training and Testing}
\label{subsec:data}
We have used corpus shared by ILSMT detailed in Table~\ref{tab:data} for the experiments. Testing for the experiments was done using Test1 which is the development set. The submitted systems were evaluated against Test2 corpus, by the organizers. For unconstrained systems, additional data~\cite{bojar:2014:monohindi,khapra:2010} has been used for language modeling.

\begin{table}[!hbt]
	\centering
	\begin{tabular}{l|c|c|c}
			& health & tourism & general \\ \hline
	training (TM) & 24K & 24K & 48K \\
	training (LM) & 24K & 24K & 48K \\
	test1 & 500 & 500 & 1000 \\
	test2 & 500 & 400 & 1000 \\
	\end{tabular}
	\caption{Corpus distribution; TM: Translation Model, LM: Language Model}
	\label{tab:data}
\end{table}

\subsection{SMT System Set Up}
\label{subsec:setup}
The baseline system was setup by using the phrase-based model~\cite{och:2003,brown:1990,marcu:2002,koehn:2003,koehn:2007} and~\newcite{koehn:2007:factored} was used for factored model. The language model was trained using KenLM~\cite{heafield:2011} toolkit with modified Kneser-Ney smoothing~\cite{chen:1996}. For factored SMT training source and target side stem were used as an alignment factor. Stemming was done using a lightweight stemmer for Hindi~\cite{ramanathan:2003:stem}. For English, we used porter stemmer~\cite{minnen:2001}.

\subsection{Evaluation Metrics}
\label{subsec:eval}
The different experimental systems were compared using, BLEU~\cite{papineni:2002}, NIST~\cite{doddington:2002}, translation error rate (TER)~\cite{snover:2006:ter}. For an MT system is to be better, higher BLEU and NIST scores with lower TER are desired.

\subsection{Pre-Processing (PP)}
\label{subsec:pp}
To tackle the morphological divergence between the source and target languages for the purpose of a better SMT system, we preprocessed the source (Bengali, Marathi, Tamil, and Telugu) for suffix separation and compound word splitting ~\cite{pimpale:2014,patel:2014} prior to training and testing. To handle the structural divergence (\enhi), we used source side reordering~\cite{patel:2013}.

\subsection{Transliteration (TR)}
\label{subsec:tr}
We developed transliteration systems using ~\newcite{durrani:2014} inbuilt with the Moses tool. We used n-best transliteration output for OOV words. These candidates were then plugged in and re-scored with the language model to get the best translation for the given source sentence.

\subsection{Language Modelling}
\label{subsec:lm}
We trained 5-gram LM using KenLM with modified Kneser-Ney smoothing. LM size was quite large (ARPA-88.7GB, binary-27.4GB), even after binarization. Training LM with the huge amount of monolingual data like ~\newcite{bojar:2014:monohindi} (approx. 10GB ) required good computing resources (60GB RAM and 200GB storage space in the working dir). Also, ~\newcite{bojar:2014:monohindi} corpus contains unwanted symbols ($<$s$>$) for KenLM which we need to remove prior training.


\section{Experiments and Results}
\label{sec:results}
Table~\ref{tab:results:cons} shows evaluation scores for various systems tried under constrained submission, that is, systems trained only on the shared data. We can see the use of the preprocessing and transliteration contributed to the improvement of around 1-5 bleu points across the language pairs. Detailed evaluation scores for unconstrained systems, that is, systems using language model built on external data, are in Table~\ref{tab:results:uncons}. Significant improvement can be observed from the table when additional data was used for language modeling.

\begin{table}[!hbt]
	\centering
	\begin{tabular}{l|c|c|c|c}
		 &  & BLEU &  NIST & TER \\ \hline
	\bnhi & S1 & 30.18 & 6.888 & 48.43 \\
		& S2 & 31.23 & 6.943 & 47.27 \\
		& S3 & 32.22 & 7.092 & 46.46 \\ \hline
	\enhi & S1 & 18.76 & 5.862 & 64.05 \\
		& S2 & 23.15 & 6.037 & 59.89 \\
		& S3 & 23.44 & 6.092 & 59.58 \\ \hline
	\mrhi & S1 & 35.74 & 7.511 & 42.75 \\
		& S2 & 40.01 & 7.805 & 39.95 \\
		& S3 & 40.21 & 7.867 & 39.62 \\ \hline
	\tahi & S1 & 16.64 & 4.742 & 64.68 \\
		& S2 & 20.17 & 5.303 & 62.65 \\
		& S3 & 20.58 & 5.391 & 62.12 \\ \hline
	\tehi & S1 & 24.88 & 6.219 & 52.18 \\
		& S2 & 28.94 & 6.532 & 49.97 \\
		& S3 & 29.86 & 6.679 & 49.19
	\end{tabular}
	\caption{Scores for different systems (CONSTRAINED); S1:BL; S2:BL+PP; S3:BL+PP+TR; BL:Baseline; TR:Transliteration}
	\label{tab:results:cons}
\end{table}

\begin{table}[!hbt]
	\centering
	\begin{tabular}{l|c|c|c|c}
		&  & BLEU &  NIST & TER \\ \hline
		\bnhi & S1 & 30.18 & 6.888 & 48.43 \\
		& S2' & 31.58 & 7.080 & 47.03 \\
		& S3' & 33.77 & 7.195 & 45.52 \\ \hline
		\enhi & S1 & 18.76 & 5.862 & 64.05 \\
		& S2' & 19.96 & 5.922 & 63.03 \\
		& S3' & 24.00 & 6.121 & 58.99 \\ \hline
		\mrhi & S1 & 35.74 & 7.511 & 42.75 \\
		& S2' & 36.80 & 7.617 & 42.35 \\
		& S3' & 41.20 & 7.935 & 39.25 \\ \hline
		\tahi & S1 & 16.64 & 4.742 & 64.68 \\
		& S2' & 16.38 & 4.677 & 64.45 \\
		& S3' & 20.38 & 5.340 & 62.25 \\ \hline
		\tehi & S1 & 24.88 & 6.219 & 52.18 \\
		& S2' & 25.47 & 6.251 & 51.99 \\
		& S3' & 29.72 & 6.669 & 49.26 \\
	\end{tabular}
	\caption{Scores for different systems (UNCONSTRAINED); S1:BL; S2':BL+ELM; S2':BL+PP+ELM; BL:Baseline; ELM:Extended LM (Additional data used for LM training)}
	\label{tab:results:uncons}
\end{table}


\section{Submission to the Shared Task}
\label{sec:submit}
We submitted two different results for the contest, namely constrained and unconstrained. The constrained systems were trained only on the data shared by the organizers. For unconstrained systems, we used additional monolingual data which include~\newcite{bojar:2014:monohindi} and~\newcite{khapra:2010}, for language modeling.

\section{Conclusion and Feature Work}
\label{sec:conclution}	
In this paper, we presented systems for translation from Bengali, English, Marathi, Tamil and Telugu to Hindi. These SMT systems with the use of source side suffix separation, compound splitting, preordering and bigger language model shows significantly higher accuracy over the baseline. Adding more complex features for factored models and formulating preprocessing with better way could be the next step.

\bibliography{acl2016}
\bibliographystyle{acl2016}

\end{document}